OpenStudy (anonymous):

I need help with a couple 9th grade algebra 1 problems. Solve: 4(x − 1) − (x + 6) = 2 Solve: 7(x − 1) = 5(x + 2) + 1 Solve: −2(x − 5) = −4 Solve: −4(x + 10) − 6 = −3(x − 2) Solve for r: C = 2^3.14r

OpenStudy (anonymous):

1) x=4 2) x=9 3) x=7 4) x=-52

OpenStudy (anonymous):

Thank you but i have one more problem if you dont mind... Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (anonymous):

It's quite simple just simplify and arrange them 1. 4(x - 1) - (x + 6) = 2 So we got to take out the parenthesis 4x - 4 - x - 6 = 2 4x - x - 4 - 6 =2 3x - 10 = 2 3x = 2 + 10 3x = 12 x = 12/3 x = 4 2. 7(x - 1) = 5(x + 2) + 1 7x - 7 = 5x + 10 + 1 7x - 7 = 5x + 11 7x - 5x = 11 + 7 2x = 18 x = 18/2 x = 9 3. -2(x - 5) = -4 -2x + 10 = -4 -2x = -14 x = -14/-2 x = 7 4. -4(x + 10) - 6 = -3(x - 2) -4x - 40 - 6 = -3x + 6 -4x - 46 = -3x + 6 -46 - 6 = -3x + 4x -52 = x x = -52 5. i don't get the problem

OpenStudy (anonymous):

for the sam problem, he is 10 right now

OpenStudy (anonymous):

he was 5 before 5 * 5 = 25 25 + 5 = 30 30 / 3 = 10 5 + 5 = 10 so it works